• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name pking

    Log entry time 00:04:08 on October 18, 2009

    Entry number 295426

    keyword=Swing shift summary

    SL : Paul King
    TO: Jeong Han Lee
    3rd: ZAFAR AHMED


    Production runs at 98 uA with IHWP OUT:

    Production runs at 15 uA with IHWP OUT:


    We had tried paging Ed twice, but the pages didn't go through. I'm pretty sure that I put in the correct number (584-7857), or maybe I made a mistake. But, for the night, if the pager doesn't work, try his cell number; it is written up on the white board.


    Shift begins with HAPPEX run 13350 in progress. Megan and Alex have taken control of the Compton DAQ.

    16:47---Start new HAPPEX run, 13351.

    17:45---End HAPPEX run, 13351. Request 15 uA for Compton coincidence measurement. Start 15 uA HAPPEX run, 13352.

    17:59---Try to check spot++; L-HRS ROC4 doesn't respond. Reboot L-HRS ROC4, restart L-HRS CODA. Take an L-HRS run (25068) to check spot++. Seems okay.

    18:11---Request 98 uA. Stop HAPPEX run 13352. Compton DAQ control is handed back to the shift crew.

    Start new HAPPEX run, 13353. Start new Compton run, 20472.

    19:09---Start new HAPPEX run, 13354.

    19:54---Start new Compton run, 20473.

    20:06---Start new HAPPEX run, 13355.

    20:33---Check spot++ with run 25069.

    21:03---Start new HAPPEX run, 13356. Several long trips in the run.

    21:12---Start new HAPPEX run, 13357. Beam trips shortly after the start and stays down.

    RF problems and a misbehaving trim card?

    22:01---OPS alerts us to an ESR alarm related to Hall A. We find the Left Q3 helium is registering as 100% full. Contact RC. Page Ed.

    22:25---We're contacted by Robert from Cryo; the high level isn't causing a problem for them, but could be a problem for us if it drops rapidly. From the Cryo Log entry 1492942:

    Left Q3 LHe level climbed to 100% around 21:30. No change in valve
    positions on the magnet seen before the level started climbing. Changed
    the Bottom Fill valve min position from -2.5 to -6 to try and compensate.

    22:30---Page Tech-on-call (Ed Folts) again.

    22:40---Ed calls in; he never got the pages. He will try to get in contact with Robert.

    22:46---Compton seems to have been locked up while I was trying to understand what was going on with the cryo. coda reboot.

    Start Compton run 20475.

    23:06---Start new HAPPEX run, 13359.

    00:03---Start new HAPPEX run, 13360.